OpenStudy (angelic):

Sagari is making a cylindrical pottery container to hold beverages. she wants the container to hold at least one quart of liquid.(1 qt.=57 3/4 cu. in., pi=22/7. if the inside diameter is three inches, how high must the inside height be

OpenStudy (mertsj):

\[57\frac{3}{4}=\frac{22}{7}(\frac{3}{2})^2h\]

OpenStudy (mertsj):

Solve for h